OpenStudy (anonymous):

please help me solve 20g^2=16g

OpenStudy (anonymous):

Well what do you think you have to do to?

OpenStudy (mathstudent55):

Subtract 16g to the left side and end up with an equation that equals zero. Then factor the left side and set each factor equal to zero. Then solve each equation.

OpenStudy (anonymous):

I'm still lost. -____-

OpenStudy (mathstudent55):

20g^2 = 16g 20g^2 - 16g = 0 Now can you factor the left side?

OpenStudy (mathstudent55):

20g^2 - 16g = 0 4g(5g - 4) = 0 4g = 0 or 5g - 4 = 0 g = 0 or 5g = 4 g = 0 or g = 4/5

OpenStudy (anonymous):

Ok heres a better way to do it. Since you want to get g by itself you simply divide by 20. Then square both sides and there is your answer. 20g^2/20=16g/20 g^2=4/5 g = rad 4/5or rad 0.80

OpenStudy (anonymous):

oh but u can simplify it more. Ill let u figure it.

OpenStudy (mathstudent55):

@blahblah_who_cares Your answer is incorrret, and your method does not work.